AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 17:18:14 2025
Author
*
:
Subject
*
:
<p>Biased detectors this morning to 150V, the leakage current for all detectors appears stable and is ~3.5 uA. We have also replaced the CAEN bias module so that all channels are now functional. Detectors are connected as follows;</p> <p>CHN0 -> Top left</p> <p>CHN1 -> Top Right</p> <p>CHN2 -> Bottom Left</p> <p>CHN3 -> Bottom Right</p> <p>Reconnected the motors and moved the detectors while biased. No observable change in the leakage current during movement at 150V. An increase in the pressure of the electron target was observed during movement (up to 1e-9 mbar). This increase was only momentarily, and the pressure recovered after a couple of seconds. We expect the pressure increase from movement to decrease over time.</p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown